Match the parametric equations to the shape of their graphs.


x = 2t + 1

y = 1 - t

shape:


x = t - t2

y = 1/2(t -3)

shape:


x = cos(t)

y = sin(t)

shape:


x = t2 - 1

y = sin(t)

shape:


1
Expert's answer
2021-04-19T16:34:14-0400

1. t=(x-1)/2

y=-x/2+1.5

shape: line

2. t=2y+3

x=2y+3-4y2-9=-4y2+2y-6

shape: parabola

3.

"x^2=cos^2t=1-sin^2t"

"y^2=sin^2t"

"x^2=1-y^2"

"x^2+y^2=1"

circle

4. "t=\sqrt{x+1}"

"y=sin(\sqrt{x+1})"

shape: sinusoid, x>-1